using Greens theorem in the plane evaluate ∮(3x+4y)dx+(2x-3y)dy where C is the circle x²+y²=4 traversed in the counterclockwise sense


1
Expert's answer
2021-12-22T03:35:24-0500
"P(x, y)=3x+4y, \dfrac{\partial P}{\partial y}=4"

"Q(x, y)=2x-y, \dfrac{\partial Q}{\partial x}=2"

Applying Green’s Theorem, you then have


"\u222e_C(3x+4y)dx+(2x-3y)dy"

"=\int \int _D(2-4)dxdy"

Transforming to polar coordinates, we obtain



"\u222e_C(3x+4y)dx+(2x-3y)dy"

"=\int \int _D(2-4)dxdy"


"=\displaystyle\int_{0}^{2\pi}\displaystyle\int_{0}^{2}(-2)rdrd\theta"

"=-2\pi(2)^2=-8\pi"